(x - 9)(x - 3) = -8 Solve by factoring and using the principle of zero products.
Answer by mananth(16946)

(x-9(x-3)=-8
x^2-3x-9x+27=-8
x^2-12x+35=0
x^2-7x-5x+35=0
x(x-7)-5(x-7)=0
(x-5)(x-7)=0
x=5 OR x=7
Ananth